OpenStudy (anonymous):

Rectangle A has an area of 16 - x2. Rectangle B has an area of x2 + 2x - 24. In simplest form, what is the ratio of the area of Rectangle A to the area of Rectangle B? Show your work.

OpenStudy (anonymous):

\[\frac{16-x^2}{x^2+2x-24}\] \[\frac{(4-x)(4+x)}{(x+6)(x-4)}\] \[-\frac{4+x}{x+6}\]

OpenStudy (anonymous):